OpenStudy (anonymous):

write quiotent as rational expression and simplify 1. 7 divided by 3/5 2. - 8/7 divided by - 4/3

OpenStudy (anonymous):

1. 7 x (5/3) = 35 / 3 2. (-8/7) x (-3/4) = 24/28 = 6/7

OpenStudy (anonymous):

Can you give me a hand with this one? 3ab/4cd divided by 9a/8d.

OpenStudy (anonymous):

Since it is division you multiply by the reciprocal. 3ab/4cd x 8d/9a 24abd/36acd Simplify. 2b/3c

OpenStudy (anonymous):

Thank you